In 1897 JJ Thomson did experiment with kinetic energies of 200 eV which passed through a pair of plates with 2 cm separation. Use a calculations to explain why JJ Thomson saw no evidence for wavelike behaviour of elecrons.

1
Expert's answer
2021-10-31T18:15:34-0400

"E=\frac{hc}{\lambda},\implies"

"\lambda=\frac{hc}E,"

"\lambda=\frac{6.63\cdot 10^{-34}\cdot 3\cdot 10^8}{200\cdot 1.6\cdot 10^{-19}}=0.62~\mu m\ll 2~cm."